OpenStudy (anonymous):

please help i will give medal solving mixed equations -2(r – 2) – 7 + 8r – r = 32

OpenStudy (anonymous):

mathstudent55 please help

OpenStudy (anonymous):

i will give you a medal and fan you

OpenStudy (mathstudent55):

The first step you need to do is to combine like terms on the left side. \(-2(r – 2) – 7 + \color{red}{8r} \color{red}{– r} = 32\) The like terms for now are in red. What is 8r - r ?

OpenStudy (mathstudent55):

Remember that r is the same as 1r. What is 8 cars - 1 car? What is 8 oranges - 1 orange? What is 8 books - 1 book? What is 8r - 1r?

OpenStudy (mathstudent55):

I thought you needed help. Are you still there?

OpenStudy (anonymous):

yah

OpenStudy (mathstudent55):

Ok. What is 8r - r ?

OpenStudy (anonymous):

5?

OpenStudy (mathstudent55):

Did you see my examples above? "Remember that r is the same as 1r. What is 8 cars - 1 car? What is 8 oranges - 1 orange? What is 8 books - 1 book? What is 8r - 1r?" You have 8 r's. You subtract one r. How many r's are you left with?

OpenStudy (mathstudent55):

How did you get 5?

OpenStudy (anonymous):

ohh 7 whoops

OpenStudy (mathstudent55):

8r - 1r = 7r, not just 7.

OpenStudy (anonymous):

ok sorry

OpenStudy (mathstudent55):

\(-2(r – 2) – 7 + \color{red}{8r} \color{red}{– r} = 32\) \(-2(r – 2) – 7 + \color{red}{7r} = 32\) We are here now. Now we need to work on the parentheses on the left side (in red below). \(\color{red}{-2(r – 2)} – 7 + 7r = 32\) Do you know the distributive property?

OpenStudy (anonymous):

not really

OpenStudy (mathstudent55):

Here is an example. \(2(x + 3) = 2 \times x + 2 \times 3 = 2x + 6\) When you have a number or variable outside parentheses multiplying a sum inside the parentheses, multiply the outside number or variable by each number or variable inside, and add.

OpenStudy (mathstudent55):